The shareholders’ equity section of the balance sheet of TNL Systems Inc. included the following accounts at December 31, 2020:

Shareholders' Equity ($ in millions)
Common stock, 270 million shares at $1 par $ 270
Paid-in capital—excess of par 1,890
Paid-in capital—share repurchase 2
Retained earnings 1,700


Required:
1. During 2021, TNL Systems reacquired shares of its common stock and later sold shares in two separate transactions. Prepare the entries for both the purchase and subsequent resale of the shares assuming the shares are (a) retired and (b) viewed as treasury stock.

  1. On February 5, 2021, TNL Systems purchased 7 million shares at $11 per share.
  2. On July 9, 2021, the corporation sold 3 million shares at $13 per share.
  3. On November 14, 2023, the corporation sold 3 million shares at $8 per share.


2. Prepare the shareholders’ equity section of TNL Systems’ balance sheet at December 31, 2023, comparing the two approaches. Assume all net income earned in 2021–2023 was distributed to shareholders as cash dividends.
Answer #1

1A. Retirement:

February 5, 2021

Common stock(7m*$1)

Paid in capital in excess of par(1890*7/270)

paid in capital share puchase

retained earnings

cash(7m*$11)

(To record purchase and retirement of common stock)

7

49

2

19

77

July 9, 2021

Cash (3m*$13)

Common stock(3*1)

Paid in capital excess of par

  (To record issuance of common stock)

39

3

36

November 14, 2023

Cash(3m*$8)

Common stock(3m*$1)

Paid in capital exess of par

(To record issuance of common stock)

24

3

21

b.Treasury stock

February 5, 2021

Treasury stock

Cash(7*$11)

(To record purchase of treasury stock)

77

77

July 9, 2021

Cash(3*13)

paid in capital share puchase

Treasury stock(3*11)

(To record sale of treasury stock)

39

6

33

November 14, 2023

Cash(3*8)

paid in capital share repuchase

Retained earnings

Treasury stock(3*11)

(To record sale of treasury stock)

24

8

1

33

2.

Shareholders equity Retirement treasury stock
Paid in capital
Common stock 269 270
Paid-in capital-excess of par 1898 1890
Retained earnings 1681 1699
Less: treasury Stock 0 11
Total Shareholder's Equity 3848 3848

Commom stock=270-7+3+3

Paid-in capital-excess of par=1890-49+36+21=1898

Retained earnings=1700-19=1681,1700-1=1699

Treasury stock=77-33-33=11